OpenStudy (anonymous):

Find the derivative of the given function. y=3x-2/2x-3

OpenStudy (mimi_x3):

Is it: \[\large \frac{3x-2}{2x-3} \] ?

OpenStudy (anonymous):

quotient rule: Lod Hi - Hi d Lo over Lo Lo

OpenStudy (anonymous):

Quotient Rule: f'(x) = g(x)f'(x) - f(x)g'(x) / (g(x))²

hero (hero):

@vi3tbloodxx , don't forget to put proper parentheses to distinguish the numerator: f'(x) = (g(x)f'(x) - f(x)g'(x)) / (g(x))²

OpenStudy (anonymous):

what mimi said

OpenStudy (anonymous):

yea that's what i meant sorry... ((3x-2)'(2x-3) - (2x-3)'(3x-2)) / (2x-3)² = (3(2x-3)-2(3x-2)) / (2x-3)² = (6x-9-6x+4) / (2x-3)² = -5/(2x-3)²

OpenStudy (mimi_x3):

or.. \[\large y' = \frac{u'v-uv'}{v^{2}} \] u = 3x-2 => u' =3 v = 2x-3 => v'=2

OpenStudy (anonymous):

i think i got it: \[3(2x-3)-2(3x-2)/(2x-3)^2\]

OpenStudy (anonymous):

is that the same as vi3tbloodxx?

OpenStudy (anonymous):

yes with an extra set of parentheses around like this (3(2x-3)-2(3x-2))

OpenStudy (anonymous):

oh then thanks just verifying.

OpenStudy (anonymous):

You can also use logarithmic differentiation ln(3x-2) - ln(2x-3) = (3/(3x-2)) - (2/(2x-3))